# Hope vs. Hate: Understanding User Interactions with LGBTQ+ News Content in Mainstream US News Media through the Lens of Hope Speech

### v1

@@ 6 Conclusions
In this paper, we present a comprehensive picture of LGBTQ+ discussions in mainstream US political discourse. Overseen by a health expert specializing in LGBTQ+ health for decades, we conduct a detailed annotation study that reveals political biases of raters are associated with how they rate supportive content for the LGBTQ+ community. Our study shows that such biases may perpetuate into models affecting marginalized voices. Our in-the-wild assessment of LGBTQ+ discussions reveal that negative comments about the community considerably outweigh positive discourse indicating a technological gap to ensure safer spaces for marginalized communities.
@@ 7 Ethics Statement
We do not collect any PII (Personally Identifiable Information) about the raters and compensate them above the minimum wage. Since content moderation can be potentially gruesome and affect the mental health of the raters [51], we maintain a small batch size (30 YouTube comments).
@@
Our substantive findings rely on fine-tuned large language models. Studies indicate that these models have a wide range of biases that reflect the texts on which they were originally trained, and which may percolate to downstream tasks [52].
@@ 8 Limitations
Our focus on YouTube comments for major news channels, while providing a rich dataset, is only a fraction of the online discource related to LGBTQ+. Platforms like Twitter, Reddit, and others have their own unique demographics, content, and user interactions. What we observed on YouTube may not be true for the rest of the internet, potentially limiting our generalizability across social media.
@@
Additionally, our study is grounded in US politics, where societal attitudes towards the LGBTQ+ have been affected by historical and legal context. However, LGBTQ+ rights and acceptance may vary dramatically across the world. In some countries, open support for LGBTQ+ rights might be more common and less contentious, while in others, it could be far more risky or even illegal to express such support. This global variation in LGBTQ+ rights and societal attitudes means that the patterns of hope speech and the very definition of what constitutes supportive language could differ significantly across cultural and national boundaries.
@@
While our study did not specifically aim at participatory AI [53, 54] as our key goal was to study the interplay of rater politics and how they perceive LGBTQ+ discussions, a considerable fraction of our raters self-reported as being part of the community. Future studies can solely focus on participatory AI involving more raters from the LGBTQ+ community and extend this research to vicarious interactions [21].
@@
We also acknowledge that many challenges faced by individual groups within the LGBTQ+ community are unique. For instance, the trans community faces several additional challenges such as participation in competitive sports or access to gender reassignment treatment. Future studies solely focusing on the trans community will add further value to prior literature focusing on specific exclusionary behavior (see, e.g., [55]).
@@
In terms of political affiliations, while we attempted to capture the broad ideological difference present in the US, this may oversimplify their nuanced reality. Our categorization of political affiliations into the groups of "Democrat," "Republican," and "Independent" may not fully capture these nuances. In reality political views exist on a spectrum rather than in discrete categories.
@@
Appendix A Additional Related Work
A.1 Additional Literature on Counter Speech
@@
There exists a large body of research done into the automatic detection of hate-speech online. This is mainly relevant in the case of social media platforms such as Facebook or YouTube, where such hateful content can be very harmful to both user’s online and in-person experiences, especially those belonging to marginalized or minority groups[56]. By identifying this hate-speech, either before or after it is posted, it can be censored or removed and the user’s that posted such content can be warned or banned. In extreme cases, entire social media platforms can be taken down by those hosting them[57]. More recent research has shown that censorship and suppression of offenders is not always the best way to deal with such situations. These approaches can often be unsuccessful in limiting the spread of hateful ideas and may even backfire, both in the context of social media and in society as a whole [7].
@@
Counter-speech has been proposed as a method of mitigating the effects of and preventing hate-speech, without the need to limit free speech through censorship. Counter-speech is a somewhat vague term and current literature defines it in a multitude of ways and acknowledges that there are many different types and strategies of counter-speech that can be found online, each achieving varying levels of success in different communities. It was found that empathy-based approaches could cause posters of xenophobic hate-speech to delete their comments and be less likely to post similar content in the near feature [58]. However, it has also been shown that strategies such as utilizing a ’Positive Tone’ or affiliating oneself with the marginalized group are often not effective and may even garner replies by other users stating that such comments will not change the opinions of the hate speakers [8]. Another negative aspect of relying on counter-speech to moderate online communities is that it requires the existence of hate-speech to be employed. By most working definitions of counter-speech, it must be in reply to a hateful comment. This limits its effectiveness as it can only be used as a response and not proactively.
A.2 Additional Literature on Hope Speech
@@
Addressing the drawbacks of counter-speech and the complexities of censorship, the emerging field of hope speech research presents a promising approach. It focuses on promoting positive, inclusive dialogues, aiming to transform and de-escalate online communities into environments of support and mutual respect. The term hope speech started to be used a few years ago by authors looking for user posted content that aimed to diffuse conflict, for example, in the context of the tensions between Indian and Pakistan [9]. Before this, similar work had been done, but used other terms such as "help-speech" to refer to the positive discourse [14].
@@
Since then, there have been multiple shared tasks in hope speech detection with over 50 teams presenting their findings [59] [60]. These shared tasks utilized datasets specifically created with hope-speech detection in mind, such as HopeEDI: a multi-lingual collection of almost 60,000 YouTube comments manually labeled as containing hope-speech or not [10]. This dataset makes a point to contain hope speech directed at LGBTQ+ communities, among many other topics. Our study is different as it focuses directly on the LGBTQ+ community, but also views it through the lens of the US political divide.
@@
One related study that didn’t explicitly focus on hope speech but provided valuable advancement to the field is the study of dehumanization. Dehumanization involves perceiving or treating people as less than human and often leads to extreme intergroup bias and hate speech. Mendelsohn, Tsvetkov, and Jurafsky developed a computational linguistic framework for analyzing dehumanizing language and applied it to discussions of LGBTQ people in the New York Times from 1986 to 2015 [5]. They found increasingly humanizing descriptions of LGBTQ people over time. Notably, different words with similar meanings, such as "gay" and "homosexual," had different levels of dehumanization, with "homosexual" being more associated with dehumanizing attitudes.
@@
While much of the previous literature on hope speech detection mentions the LGBTQ+, as the task gravitates towards marginalized groups, there are gaps in research focusing directly on it. García-Baena et al. created the dataset SpanishHopeEDI. It contains 1,650 annotated Spanish LGBTQ+-related tweets. A multitude of traditional machine-learning algorithms were tested on identifying hope-speech on them, leading to very impressive results [61]. One related shared task similar to the previous ones was conducted focusing on this topic, however it was for hate-speech detection instead. For this, a dataset that specifically labelled LGBTQ+-related comments as homophobic or transphobic was used [62].
@@
Another important question for hope speech detection is understanding how subjective interpretations of language can impact the effectiveness of content moderation. A related study investigates the disagreement between human and machine moderators on what constitutes offensive speech in political discourse [21]. This research highlights the challenges of subjectivity in moderation, revealing that both human and machine moderators often disagree on what is considered offensive, especially when political leanings are involved. These findings are relevant for hope speech detection as they underscore the need for nuanced understanding and handling of subjective content, ensuring that genuine hope speech is accurately identified amidst varying interpretations. In the same way that there can be disagreement on what is offensive, there can be disagreement on what is supportive or hopeful.
@@
The recent introduction and consistent improvement of LLMs in the field of natural language processing has opened many new doors, and the task of hope speech detection is no different. However, at this point in time there has not been a lot of research into how effective LLMs are at hope-speech detection, never mind LGBTQ+ hope speech superficially. At the time of writing, only one paper has looked into this and attempted zero-shot ChatGPT prompting to label a Hope Speech dataset [63]. They found that it resulted in poor performance for English texts, but was very successful for Spanish text. More work has been done surrounding LLM performance into hate-speech detection. The first extensive study into this found that prompting strategies had a significant impact on results [64]. They found using a Chain-of-Thought Reasoning Prompt had the highest results in all metrics and overall was effective in detecting hate-speech.
Appendix B Prompts Used
@@
Appendix C Annotation Guidelines
@@
Appendix D Two-Label Agreement
@@
Table 12: Two-Label Political Affiliation Agreement (Human Annotators) Values are Fleiss’ $\kappa$
@@
Table 13: Two-Label LGBTQ+ Agreement (Human Annotators) Values are Fleiss’ $\kappa$
Appendix E Two-Label Classification Prompt
@@
Appendix F Curating $\mathcal{D}_{\textit{hope}}$
F.0.1 Initial Human and LLM Annotation
Initially, one of the authors label a set of 1,500 comments (500 randomly sampled comments from three news outlets hosting $\mathcal{V}_{\textit{LGBTQ+}}$ videos) as being hope speech directed towards the LGBTQ+ community or not. See Appendix A for the guidelines used in this process. This step only identified 80 positives (5.33%) i.e., LGBTQ+ hope speech suggesting that hope speech is a rare class. In order to curate a balanced dataset in a time- and cost-effective way, we turn to leveraging human-LLM collaborative annotation in the following way [65]. Using the initial labelled set of 1,500 comments, we fine-tune Llama2-13B (training details are in the Appendix). While this fine-tuned classifier was not highly accurate, it allowed us to identify hope speech more efficiently. We use this model to classify sets of 20,000 unseen comments from each channel. The same author as before went through the positives and verified any true positives until we obtain a total of 975 verified LBGTQ+ hope speech comments; 325 from each channel. These comments combined with a set of 975 random negatives labeled by the model, balanced for each channel, made up the 1,950 comment corpus that would be labelled by our crowd-sourced annotators as our seed set.
F.0.2 Active Learning
The first was minority certainty sampling. Prior literature shows that this method can be useful for reducing class imbalance [14, 66]. We consider our best-performing model at this point (a fine-tuned Llama) and have it classify an unseen set of 50k randomly sampled comments from each channel. We took the 150 comments of the highest certainty with labels Negative and Neutral, for each channel, totaling 900 instances. We used the same crowd-sourced annotation design described in Section 3.3.1 to have these labelled. This allowed us to help close the gap in imbalance for these two labels. Our final dataset at this step was 2,850 instances. Looking at the distribution of comments with consensus among the annotators, we had 680 Positives, 682 Negatives, 227 Neutrals, and 764 Irrelevants. We found that this improved our model’s performance, especially with the Neutral class, and were interested in seeing if continue sampling could improve it more.
Appendix G Homophobic Survey Feedback Appendix
Survey Feedback 1:   You can identify as anything you want However how you were born cannot be changed or altered. Surgery and clothing does not change what you were born as. Time for common sense to reign again. 1,000 years from now when somebody digs up bones you will never hear them say oh look we just found a tranny!!!!
Survey Feedback 2:   Comments speaking truth about the alphabet mafia are not negative just because the alphabet mafia does not like being confronted with truth.
Appendix H Licenses
Appendix I Packages
Package Version
PyTorch 2.3.0-rc12
Transformers 4.35.2
Datasets 2.8.0
NumPy 1.26.3
Scikit-learn 1.4.0
PEFT 0.5.0
Table 14: Software packages and versions
Appendix J Computational Resources
Appendix K Training Setup and Hyperparameters
Hyperparameter Value
Learning Rate 2e-4
Batch Size 8
Number of Epochs 5
Weight Decay 0.01
LoRA rank (r) 8
LoRA alpha 32
LoRA dropout 0.1
Table 15: Hyperparameters used during our fine-tuning
Appendix L Annotator Demographics

### v2

@@ 6 Conclusions
This paper presents a comprehensive picture of LGBTQ+ discussions in mainstream US political discourse. Overseen by a health expert specializing in LGBTQ+ health for decades, we conduct a detailed annotation study that reveals political biases of raters are associated with how they rate supportive content for the LGBTQ+ community. Our study shows that such biases may perpetuate into models affecting marginalized voices. Our in-the-wild assessment of LGBTQ+ discussions reveal that negative comments about the community considerably outweigh positive discourse indicating a technological gap to ensure safer spaces for marginalized communities.
@@ 7 Ethics Statement
We do not collect any PII (Personally Identifiable Information) about the raters and compensate them above the minimum wage. Since content moderation can be potentially gruesome and affect the mental health of the raters Solon (2017), we maintain a small batch size (30 YouTube comments).
@@
Our substantive findings rely on fine-tuned LLM. Studies indicate that these models have a wide range of biases that reflect the texts on which they were originally trained, and which may percolate to downstream tasks Bender et al. (2021).
@@ 8 Limitations
Our focus on YouTube comments for major news channels, while providing a rich dataset, is only a fraction of the online discource related to LGBTQ+. Platforms like Twitter, Reddit, and others have their own unique demographics, content, and user interactions. What we observed on YouTube may not be true for the rest of the Internet, potentially limiting our generalizability across social media. We conducted an exploratory investigation on Reddit, sampling 15,000 comments on LGBTQ+-relevant posts (2016-2024) from r/politics, r/republican, and r/democrat, for a total of 45,000 comments. Overall, the positivity ratio remains below 0.5 (0.40), meaning that negative comments outnumber hope speech on Reddit, as on YouTube. However the subreddit-level breakdown is more nuanced, the r/politics and r/democrat subreddits show a slight tilt toward hope speech, whereas r/republican is strongly negative. This differs from YouTube, where even communities one might expect to have more positivity, such as the CNN comment section, still had negative comments outnumbering hope speech. These results, seen in Table A20, underline the fact that future work across additional communities and media is necessary.
@@
Additionally, our study is grounded in US politics, where societal attitudes towards the LGBTQ+ have been affected by historical and legal context. However, LGBTQ+ rights and acceptance may vary dramatically across the world. In some countries, open support for LGBTQ+ rights might be more common and less contentious, while in others, it could be far more risky or even illegal to express such support. This global variation in LGBTQ+ rights and societal attitudes means that the patterns of hope speech and the very definition of what constitutes supportive language could differ significantly across cultural and national boundaries. Recent research Samir et al. (2024) demonstrates these cross-cultural differences concretely, finding significant variations in how LGBTQ+ individuals are portrayed across different language Wikipedias, with some versions systematically emphasizing negative biographical information or omitting positive achievements based on local cultural attitudes.
@@
While our study did not specifically aim at participatory AI Harrington et al. (2019); Birhane et al. (2022) as our key goal was to study the interplay of rater politics and how they perceive LGBTQ+ discussions, a considerable fraction of our raters self-reported as being part of the community. Future studies can solely focus on participatory AI involving more raters from the LGBTQ+ community and extend this research to vicarious interactions Weerasooriya et al. (2023).
@@
We also acknowledge that many challenges faced by individual groups within the LGBTQ+ community are unique. For instance, the trans community faces several additional challenges such as participation in competitive sports or access to gender reassignment treatment. Future studies solely focusing on the trans community will add further value to prior literature focusing on specific exclusionary behavior (see, e.g., Lu and Jurgens, 2022).
@@
In terms of political affiliations, while we attempted to capture the broad ideological difference present in the US, this may oversimplify their nuanced reality. Our categorization of political affiliations into the groups of "Democrat," "Republican," and "Independent" may not fully capture these nuances. In reality, political views exist on a spectrum rather than in discrete categories.
While we have attempted to make the definition of hope speech clear, operationalizable, and indicate how it differs from positive sentiment, we acknowledge room for further nuance. Appendix E lists a nuanced definition of hope speech which we considered in our initial annotation phase conducted by the author. However, our initial pilot revealed that this fine-grained definition resulted in poor quality annotation and low annotator agreement. Keeping rater instructions simple to reduce cognitive load on raters is a recommended practice in the crowd-sourcing literature Finnerty et al. (2013). To reduce cognitive load, we adopted a simpler four-way scheme, Positive, Negative, Neutral, Irrelevant, and worded our annotation guidelines as advised by the subject-matter expert using the more familiar term Positive in place of Hope Speech. Although this worked well for the purposes of our study, a finer-grained taxonomy and clearer operational guidelines merit deeper investigation.
@@
9 Acknowledgements
We acknowledge Research Computing at the Rochester Institute of Technology for providing computational resources. We thank Pranita Menavlikar and Pavan Turlapati for conducting preliminary investigations that informed this work, and Logan Crowl for assisting with the implementation of the crowdsourced experiments, along with the human annotators who took part in the study. We also thank Dr. Naveen Sharma and Dr. Christian Newman for their support and feedback as part of Jonathan Pofcher’s thesis committee. KhudaBukhsh were partly supported by a gift from Lenovo.
Appendix A Curating $\mathcal{D}_{\textit{hope}}$
A.1 Initial Human and LLM Annotation
Initially, one of the authors label a set of 1,500 comments (500 randomly sampled comments from three news outlets hosting $\mathcal{V}_{\textit{LGBTQ+}}$ videos) as being hope speech directed towards the LGBTQ+ community or not. See Appendix A for the guidelines used in this process. This step only identified 80 positives (5.33%) i.e., LGBTQ+ hope speech suggesting that hope speech is a rare class. In order to curate a balanced dataset in a time- and cost-effective way, we turn to leveraging human-LLM collaborative annotation in the following way Wang et al. (2024). Using the initial labelled set of 1,500 comments, we fine-tune Llama2-13B (training details are in the Appendix). While this fine-tuned classifier was not highly accurate, it allowed us to identify hope speech more efficiently. We use this model to classify sets of 20,000 unseen comments from each channel. The same author as before went through the positives and verified any true positives until we obtain a total of 975 verified LBGTQ+ hope speech comments; 325 from each channel. These comments combined with a set of 975 random negatives labeled by the model, balanced for each channel, made up the 1,950 comment corpus that would be labelled by our crowd-sourced annotators as our seed set.
A.2 Active Learning
The first was minority certainty sampling. Prior literature shows that this method can be useful for reducing class imbalance Palakodety et al. (2020a); Attenberg et al. (2010). We consider our best-performing model at this point (a fine-tuned Llama) and have it classify an unseen set of 50k randomly sampled comments from each channel. We took the 150 comments of the highest certainty with labels Negative and Neutral, for each channel, totaling 900 instances. We used the same crowd-sourced annotation design described in Section 3.3.1 to have these labelled. This allowed us to help close the gap in imbalance for these two labels. Our final dataset at this step was 2,850 instances. Looking at the distribution of comments with consensus among the annotators, we had 680 Positives, 682 Negatives, 227 Neutrals, and 764 Irrelevants. We found that this improved our model’s performance, especially with the Neutral class, and were interested in seeing if continue sampling could improve it more.
Appendix B Comparison to Off-The-Shelf Classifiers
Table A1 compares the performance of our hope speech classifier against traditional baselines for sentiment classification, hate speech detection, and stance detection.
Classifier Macro F1 Pos. F1 Neg. F1
Traditional Sentiment (DistilBERT-sst2) 0.578 0.402
Traditional Sentiment (Positive/Negative only) 0.601 0.52 0.682
Toxicity-based (Detoxify-toxic-bert) 0.503 0.129
Hate Speech (cardiffnlp hate-roBERTa) 0.605 0.258
GPT-4o (Classification Prompt) 0.666 0.862 0.590
Our best model (Fine-tuned Llama) 0.716 0.866 0.684
Table A1: Performance comparison for off-the-shelf classifiers on relevant tasks. For classifiers optimized to detect supportive content, the Positive F1 is reported on a binary formulation mapping hope speech (Positive) versus non-hope speech (Negative); for toxicity and hate speech detectors the Negative F1 is shown on a binary formulation mapping hate speech (Negative) versus non-hate speech (Positive, Neutral, Irrelevant). Blank cells indicate metrics that do not apply.
Appendix C Additional Related Work
C.1 Additional Literature on Counter Speech
@@
There exists a large body of research done into the automatic detection of hate-speech online. This is mainly relevant in the case of social media platforms such as Facebook or YouTube, where such hateful content can be very harmful to both user’s online and in-person experiences, especially those belonging to marginalized or minority groupsMüller and Schwarz (2020). By identifying this hate-speech, either before or after it is posted, it can be censored or removed and the user’s that posted such content can be warned or banned. In extreme cases, entire social media platforms can be taken down by those hosting themAgarwal et al. (2022). More recent research has shown that censorship and suppression of offenders is not always the best way to deal with such situations. These approaches can often be unsuccessful in limiting the spread of hateful ideas and may even backfire, both in the context of social media and in society as a whole Benesch (2014).
@@
Counter-speech has been proposed as a method of mitigating the effects of and preventing hate-speech, without the need to limit free speech through censorship. Counter-speech is a somewhat vague term and current literature defines it in a multitude of ways and acknowledges that there are many different types and strategies of counter-speech that can be found online, each achieving varying levels of success in different communities. It was found that empathy-based approaches could cause posters of xenophobic hate-speech to delete their comments and be less likely to post similar content in the near feature Hangartner et al. (2021). However, it has also been shown that strategies such as utilizing a ’Positive Tone’ or affiliating oneself with the marginalized group are often not effective and may even garner replies by other users stating that such comments will not change the opinions of the hate speakers Mathew et al. (2019). Another negative aspect of relying on counter-speech to moderate online communities is that it requires the existence of hate-speech to be employed. By most working definitions of counter-speech, it must be in reply to a hateful comment. This limits its effectiveness as it can only be used as a response and not proactively.
C.2 Additional Literature on Hope Speech
@@
Addressing the drawbacks of counter-speech and the complexities of censorship, the emerging field of hope speech research presents a promising approach. It focuses on promoting positive, inclusive dialogues, aiming to transform and de-escalate online communities into environments of support and mutual respect. The term hope speech started to be used a few years ago by authors looking for user posted content that aimed to diffuse conflict, for example, in the context of the tensions between Indian and Pakistan Palakodety et al. (2020b). Before this, similar work had been done, but used other terms such as "help-speech" to refer to the positive discourse Palakodety et al. (2020a).
@@
Since then, there have been multiple shared tasks in hope speech detection with over 50 teams presenting their findings Chakravarthi et al. (2022a) Chakravarthi and Muralidaran (2021). These shared tasks utilized datasets specifically created with hope-speech detection in mind, such as HopeEDI: a multi-lingual collection of almost 60,000 YouTube comments manually labeled as containing hope-speech or not Chakravarthi (2020). This dataset makes a point to contain hope speech directed at LGBTQ+ communities, among many other topics. Our study is different as it focuses directly on the LGBTQ+ community, but also views it through the lens of the US political divide.
@@
One related study that didn’t explicitly focus on hope speech but provided valuable advancement to the field is the study of dehumanization. Dehumanization involves perceiving or treating people as less than human and often leads to extreme intergroup bias and hate speech. Mendelsohn, Tsvetkov, and Jurafsky developed a computational linguistic framework for analyzing dehumanizing language and applied it to discussions of LGBTQ people in the New York Times from 1986 to 2015 Mendelsohn et al. (2020). They found increasingly humanizing descriptions of LGBTQ people over time. Notably, different words with similar meanings, such as "gay" and "homosexual," had different levels of dehumanization, with "homosexual" being more associated with dehumanizing attitudes.
@@
While much of the previous literature on hope speech detection mentions the LGBTQ+, as the task gravitates towards marginalized groups, there are gaps in research focusing directly on it. García‑Baena et al. created the dataset SpanishHopeEDI. It contains 1,650 annotated Spanish LGBTQ+-related tweets. A multitude of traditional machine-learning algorithms were tested on identifying hope-speech on them, leading to very impressive results García-Baena et al. (2023). One related shared task similar to the previous ones was conducted focusing on this topic, however it was for hate-speech detection instead. For this, a dataset that specifically labelled LGBTQ+-related comments as homophobic or transphobic was used Chakravarthi et al. (2022b).
@@
Another important question for hope speech detection is understanding how subjective interpretations of language can impact the effectiveness of content moderation. A related study investigates the disagreement between human and machine moderators on what constitutes offensive speech in political discourse Weerasooriya et al. (2023). This research highlights the challenges of subjectivity in moderation, revealing that both human and machine moderators often disagree on what is considered offensive, especially when political leanings are involved. These findings are relevant for hope speech detection as they underscore the need for nuanced understanding and handling of subjective content, ensuring that genuine hope speech is accurately identified amidst varying interpretations. In the same way that there can be disagreement on what is offensive, there can be disagreement on what is supportive or hopeful.
@@
The recent introduction and consistent improvement of LLMs in the field of natural language processing has opened many new doors, and the task of hope speech detection is no different. However, at this point in time there has not been a lot of research into how effective LLMs are at hope-speech detection, never mind LGBTQ+ hope speech superficially. At the time of writing, only one paper has looked into this and attempted zero-shot ChatGPT prompting to label a Hope Speech dataset Ngo and Tran (2023). They found that it resulted in poor performance for English texts, but was very successful for Spanish text. More work has been done surrounding LLM performance into hate-speech detection. The first extensive study into this found that prompting strategies had a significant impact on results Guo et al. (2024). They found using a Chain-of-Thought Reasoning Prompt had the highest results in all metrics and overall was effective in detecting hate-speech.
Appendix D Prompts Used
@@
Appendix E Annotation Guidelines
@@
Appendix F Two-Label Agreement
@@
Table A2: Two-Label Political Affiliation Agreement (Human Annotators) Values are Fleiss’ $\kappa$
@@
Table A3: Two-Label LGBTQ+ Agreement (Human Annotators) Values are Fleiss’ $\kappa$
Appendix G Two-Label Classification Prompt
The two-label classification prompt is presented in Figure 5.
@@ Appendix G Two-Label Classification Prompt
Appendix H Crowd-sourced Study Compensation
We compensate raters \$0.10 for each instance. This yields a compensation of \$3 for each 30 comment batch. We observe a median task completion time of 11:53 minutes, implying an hourly wage estimate of \$15.16, more than the minimum wage. Rater compensation is grounded in prior literature Leonardelli et al. (2021); Bai et al. (2021); Weerasooriya et al. (2023). No participants complained about the compensation in their feedback.
Appendix I Survey Feedback
Table A4 lists a few examples of homophobic survey feedback. Table A5 lists a few examples of survey feedback expressing a positive stance towards our research task or the LGBTQ+ community.
Examples of Negative Survey Feedback
“You can identify as anything you want However how you were born cannot be changed or altered. Surgery and clothing does not change what you were born as. Time for common sense to reign again. 1,000 years from now when somebody digs up bones you will never hear them say oh look we just found a trnny!!!!”
“Comments speaking truth about the alphabet mafia are not negative just because the alphabet mafia does not like being confronted with truth.”
Table A4: Examples of survey feedback expressing homophobic or negative views.
Examples of Positive Survey Feedback
“Reading so much hateful, ignorant rhetoric was tough! Thanks for all you’re doing to spread peace and love in the world.”
“This study was very eye opening. Thank you for letting me participate in this wonderful study. Have a wonderful day.”
Table A5: Examples of survey feedback expressing support for the research and LGBTQ+ community.
Appendix J Annotator Demographics